OpenStudy (anonymous):

Solve by Factoring: 2x2 + 5x - 3 = 0 What are the solutions for x?

OpenStudy (anonymous):

(2x-1)(x+3)=0

OpenStudy (anonymous):

what are the values of x?

OpenStudy (anonymous):

these are the answers to choose from x = 2 x = 5 x = 1/2 x = -3

OpenStudy (anonymous):

what is your choice based on the factors above?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

(2x-1)(x+3)=0 for this to be true you have to have: 2x-1=0 or x+3=0 so, what are the value of x?

OpenStudy (anonymous):

In order to understand these, you would have to know what ab = 0 means

OpenStudy (anonymous):

x=1/2 x=-3 got to go :)

OpenStudy (anonymous):

when you see two terms multiplied by each other and set equal to zero, it means that either one of the two could equal zero. So if ab = 0... a = 0 or b = 0